OpenStudy (scarlettfarra2000):

1. Find the roots of the polynomial equation x^2-14x-32=0 2. Find the zeros of the polynomials function F(x) = x^2+12x+20

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle x^2-14x-32=0 }\) \(\color{#000000}{ \displaystyle x^2-14x=32 }\) \(\color{#000000}{ \displaystyle x^2-14x+49=32+49 }\) \(\color{#000000}{ \displaystyle (x-7)^2=81 }\) this is one way to do it.